inverse functions
find f(g(x)) and g(f(x)) and determine whether the functions f and g are inverse of each other.
f(x)= 6x-3 and g(x)= x+6/3

Answer by drk(1908)

F(g(x)) means put the g(x) function into the f(x) function. We get:
f(g(x)) = 6(x+6/3) - 3 = 6x +12 - 3 = 6x + 9
g(f(x)) = 6x-3 + 6/3 = 6x - 1
The functions are not inverses